# Self-Foveate: Enhancing Diversity and Difficulty of Synthesized Instructions from Unsupervised Text via Multi-Level Foveation

### v1

@@ 3.2 Scatter-foveate Level
Complete unsupervised texts often imply relationships between different entities or attributes within the text. However, instruction sets directly synthesized by teacher large language models (LLMs) may struggle to detect connections that require reasoning or deeper understanding in the text. This could result in the instruction sets synthesized by the teacher model missing deep-level information from the text, thereby affecting downstream model performance. To address this, we employ the scatter-foveate mechanism and direct synthesis method to ensure that these deep-level insights are reflected in the synthesized instruction sets. Attention to the deeper, less obvious implicit information enhances the diversity of the synthesized instruction sets, while insights into the relationships between entities or attributes increase the difficulty of the instructions.
@@
This mechanism aims to guide the teacher LLM to synthesize instructions from a deeper perspective of the unsupervised text. We extract foveate elements more broadly from the text and randomly combine them into a certain number of foveate groups based on an empirical distribution.
@@
Based on the formed foveate groups, we synthesize instructions using the direct synthesis method. The direct synthesis method treats each element in the foveate group as an indispensable part of the instruction to be synthesized, actively guiding the teacher LLM to consider the deep-level connections between different entities or attributes and solidifying such reasoning-based or deep-thinking connections in the synthesized instructions. After the instructions are synthesized, the answers are regenerated based on the text.
@@ 3.3 Macro-foveate Level
Unsupervised texts often contain content requiring focused attention, such as figurative devices and rhetorical exaggerations. Notably, teacher large language models (LLMs) may require additional prompting to specifically emphasize the critical information embedded within these writing techniques. To capture information more comprehensively from a global perspective, we employ the macro-foveate mechanism and transcription synthesis method to guide teacher LLMs in developing a profound understanding of unsupervised texts, ensuring effective extraction of key information conveyed through literary devices.
@@
This mechanism aims to guide teacher LLMs in comprehending unsupervised texts from a holistic perspective. We specifically highlight and extract text segments employing writing techniques – including metaphor, hyperbole, contrastive foil, rhetorical questioning, and citation – which are extracted as foveate segments.
@@
Based on the identified writing technique segments, we employ transcription synthesis to convert each foveate segment into an instructional format. This process transforms declarative foveate segments into interrogative or imperative forms. Subsequently, corresponding answers are synthesized according to the content of the unsupervised text.
@@ 3.4 Re-synthesis Module
Due to limitations such as the capabilities of the teacher LLM, not all instructions in a directly synthesized instruction set may be answerable by the referenced unsupervised text. To ensure the synthesized instructions remain faithful to the referenced unsupervised text, we employ a re-synthesis module to repeatedly re-synthesize these failed instructions over multiple iterations, aiming to generate instructions that can be fully answered based on the referenced unsupervised text.
Single Sample Reference-synthesis
During re-synthesis, we process only one failed instruction at a time and randomly sample a subset of successfully synthesized instructions as reference examples for the teacher LLM in each iteration. By providing the teacher LLM with different batches of reference samples during each re-synthesis attempt, we enhance the success rate of re-synthesis.
@@
The outputs of large language models (LLMs) are significantly influenced by hyperparameters such as temperature, top-p, and frequency penalty. To improve re-synthesis success rates, we meticulously calibrated these hyperparameters to define a high-creativity mode, enabling the teacher LLM to synthesize instructions more creatively based on reference samples.
Dataset Source # Test Examples
Question Context
SQuAD Crowdsourced Wikipedia 11639
HotpotQA Crowdsourced Wikipedia 2500
FilmWiki LLM Wikipedia 7398
@@
Datasets Methods Diversity Metrics
SelfBLEU Div. Embedding Div.
SQuAD Self-QA 0.593 0.838
Bonito 0.494 0.838
Wiki2023 0.550 0.842
Self-Foveate 0.665 0.851
Test Questions 0.695 0.840
HotpotQA Self-QA 0.463 0.823
Bonito 0.371 0.769
Wiki2023 0.554 0.822
Self-Foveate 0.607 0.835
Test Questions 0.634 0.786
FilmWiki Self-QA 0.406 0.687
Bonito 0.197 0.677
Wiki2023 0.341 0.664
Self-Foveate 0.563 0.706
Test Questions 0.316 0.618
@@
Model Settings GPT-4o mini DeepSeek-V3
SQuAD HotpotQA FilmWiki SQuAD HotpotQA FilmWiki
Rec. Acc. Rec. Acc. Rec. Acc. Rec. Acc. Rec. Acc. Rec. Acc.
Llama-3.1-8B None* 0.309 0.202 0.244 0.160 0.212 0.082 0.309 0.202 0.244 0.160 0.212 0.082
Self-QA 0.367 0.384 0.372 0.358 0.328 0.201 0.389 0.412 0.399 0.378 0.370 0.239
Wiki2023 0.327 0.361 0.338 0.322 0.333 0.235 0.342 0.370 0.340 0.328 0.349 0.244
Bonito* 0.386 0.405 0.360 0.372 0.219 0.153 0.386 0.405 0.360 0.372 0.219 0.153
Self-Foveate 0.484 0.490 0.507 0.486 0.512 0.367 0.481 0.491 0.525 0.501 0.548 0.397
Qwen2.5-7B None* 0.251 0.300 0.266 0.234 0.139 0.032 0.251 0.300 0.266 0.234 0.139 0.032
Self-QA 0.249 0.232 0.276 0.246 0.206 0.082 0.119 0.125 0.102 0.106 0.111 0.056
Wiki2023 0.215 0.221 0.135 0.112 0.192 0.093 0.170 0.083 0.197 0.203 0.202 0.136
Bonito* 0.143 0.109 0.212 0.199 0.168 0.098 0.143 0.109 0.212 0.199 0.168 0.098
Self-Foveate 0.408 0.414 0.372 0.329 0.283 0.140 0.388 0.389 0.342 0.331 0.261 0.140
Gemma-2-9B None* 0.224 0.121 0.175 0.078 0.211 0.099 0.224 0.121 0.175 0.078 0.221 0.099
Self-QA 0.383 0.409 0.408 0.389 0.429 0.315 0.402 0.435 0.424 0.408 0.509 0.386
Wiki2023 0.336 0.378 0.361 0.352 0.478 0.384 0.364 0.399 0.373 0.365 0.494 0.401
Bonito* 0.411 0.457 0.366 0.373 0.255 0.196 0.411 0.457 0.366 0.373 0.255 0.196
Self-Foveate 0.507 0.525 0.537 0.520 0.672 0.528 0.499 0.514 0.552 0.525 0.697 0.581
@@ 4 Experiment
In this section, our experiments focus on three critical research questions: (Q1) How effective is Self-Foveate in enhancing instruction diversity? (Q2) How effective is Self-Foveate in enhancing instruction difficulty? (Q3) How effective is Self-Foveate in improving the model’s problem-solving capabilities during instruction fine-tuning?
@@ 4.1 Experimental Setup
We employ three independent datasets, including the training set of the FilmWiki dataset containing 2,385 unsupervised texts with corresponding question-answer pairs. Additionally, we sample two widely-used benchmark QA datasets from the MRQA 2019 shared task Fisch et al. (2019): SQuAD (Rajpurkar et al., 2016) (following Bonito (Nayak et al., 2024a)) and HotpotQA (Yang et al., 2018). To maintain comparable data scale and computational costs with the FilmWiki dataset, we extract 2,500 unsupervised texts with corresponding QA pairs from each dataset’s training split as substitutes for the complete collections. Further implementation details are provided in Table 2.
@@
We consider three key baselines: zero-shot, Self-QA, Bonito and Wiki2023. For the zero-shot baseline, we prompt models for evaluation without leveraging any unsupervised texts from the target task (None). The Self-QA baseline employs an unsupervised knowledge-guided method for extracting instruction-question-answer triples (Self-QA(Zhang and Yang, 2023). The Bonito baseline utilizes a 7B-sized specialized model to generate various types of questions from unsupervised text. We configure it to produce questions of the “question answering without choices” type and obtain a sufficient number of samples through multiple sampling iterations (Bonito). The Wiki2023 baseline implements a text-based QA pair extraction methodology (Wiki2023). All baselines use identical unsupervised texts as our method, as specified in Section 4.1.
@@
As described in Section 3, we process unsupervised texts through Self-Foveate using GPT-4o-mini and DeepSeek-V3 to generate instructional data. We implement two distinct hyperparameter configurations emphasizing high stability and high creativity, respectively, with detailed specifications in Appendix B.
@@
We select three state-of-the-art open-source foundation models as our base models prior to instruction tuning: Meta-Llama-3.1-8B (Meta, 2024), Qwen2.5-7B (Team, 2024), and Gemma-2-9B. These decoder-only language models employ next-word prediction objectives and were pretrained on trillions of tokens without any instruction-based fine-tuning.
4.2 Diversity Discussion
To explicitly evaluate the diversity of the instructions in a metric-driven manner, we follow established practices recommended in Zhu et al. (2018); Perez et al. (2022); Tevet and Berant (2021) and employ two metrics: SelfBLEU score and Sentence-BERT embedding distances (Reimers and Gurevych, 2019a). These metrics capture different facets of diversity. SelfBLEU measures diversity in the form of text, while embedding distances measure diversity in the semantics of text. For SelfBLEU scores, we compute the average SelfBLEU scores using $n$-grams for $n\in\{2,3,4,5\}$, following the approach suggested by Zhu et al. (2018). We use the implementation of the SelfBLEU metric in Alihosseini et al. (2019). Further details are available in Appendix C.
Table 3 demonstrates that Self-Foveate-generated instructions achieve substantial diversity improvements in both textual and semantic dimensions, attaining or even surpassing the diversity level of crowdsourced test questions through a low-cost automated process.
4.3 Difficulty Discussion
To investigate the difficulty of the generated instructions, we designed prompts and employed GPT-4o to conduct a rigorous head-to-head comparison between the instructions generated by Self-Foveate and those generated by baseline methods. Specifically, for each unsupervised text in each data subset, we used two contrasting methods to generate an equal number of instructions. We then provided GPT-4o with a set of instructions generated by both methods under the same unsupervised text, ensuring that the instruction sets were anonymous and their relative positions were randomized. Subsequently, we recorded GPT-4o’s judgments and calculated the win rate for Self-Foveate. The results of the head-to-head comparison, as shown in Table 5, demonstrate the outstanding difficulty of the instructions synthesized by Self-Foveate. Further details can be found in Appendix D.
Figure 3: Impact of instruction set scale from Self-Foveate and baselines on model fine-tuning performance.
Dataset Baseline  Self-Foveate WR.
Win Lose
SQuAD Self-QA 70.64% 29.36%
Wiki2023 80.83% 19.17%
Bonito 99.96% 00.04%
HotpotQA Self-QA 89.52% 10.48%
Wiki2023 91.17% 08.83%
Bonito 100.00% 00.00%
FilmWiki Self-QA 85.12% 14.88%
Wiki2023 95.08% 04.92%
Bonito 96.31% 03.69%
@@ 4.4 Problem-solving Capabilities
To evaluate the impact of instruction-tuning datasets on model performance in downstream tasks, we fine-tuned open-source models and assessed their performance on in-distribution problems. Table 4 illustrates the effects of instruction sets synthesized by two models (GPT-4o mini and DeepSeek-V3) during fine-tuning on open-source model performance, with equal quantities of instructions per unsupervised text. As answers in test problems often consist of phrase-level or sentence-fragment extractions from source texts while large language models (LLMs) tend to generate more comprehensive responses, we selected both recall and LLM-evaluated accuracy as evaluation metrics. Here, higher recall indicates a greater probability of generated answers containing correct responses, while higher LLM-evaluated accuracy reflects stronger consistency between answers generated by DeepSeek-V3 and ground truth labels. To control potential confounding effects from the number of instructions synthesized per unsupervised text and examine the agreement between these two metrics, we investigated the influence of instruction set scale synthesized by Self-Foveate and baseline methods on model fine-tuning performance using the HotpotQA dataset, as shown in Figure 3. Additional details regarding downstream task evaluations are provided in Appendix E.
Table 4 demonstrates that two state-of-the-art (SOTA) models achieve optimal performance across all downstream tasks in both metrics when using instructions synthesized via Self-Foveate. Notably, during the fine-tuning process of the Qwen2.5-7B model, we observed that instruction data synthesized by baseline methods may lead to performance degradation compared to the pre-fine-tuned model, further demonstrating the advancement of the Self-Foveate method in instruction synthesis. Our analysis presented in Figure 3 reveals that as the number of instructions synthesized per unsupervised text by Self-Foveate increases, the performance of downstream fine-tuning tasks improves significantly, with an expanding performance gap compared to other baseline methods. Figure 3 also visually demonstrates the consistency between recall and LLM-evaluated accuracy, validating the appropriateness of these two metric selections.

### v2

@@ 3.2 Scatter-foveate Level
Beyond individual entities, unsupervised text often encodes implicit relationships (causal links, comparisons, or temporal dependencies) between entities or attributes that are scattered across different passages. A teacher LLM synthesizing instructions without targeted guidance is unlikely to surface these cross-entity connections, producing instruction sets that remain shallow. The scatter-foveate level tackles this by combining dispersed key information through a grouping mechanism and applying direct synthesis, so that the resulting instructions require reasoning about inter-entity relationships. This design broadens diversity by surfacing less obvious implicit information and raises difficulty by demanding cross-entity reasoning.
@@
This mechanism encourages the teacher LLM to reason about connections across different parts of the unsupervised text. We extract foveate elements more broadly from the text and randomly combine them into foveate groups, with group sizes sampled from an empirical distribution.
@@
Based on the formed foveate groups, we synthesize instructions using the direct synthesis method, which treats each element in the foveate group as an indispensable component of the target instruction, guiding the teacher LLM to uncover implicit semantic connections between different entities or attributes and to embed such cross-entity reasoning in the synthesized instructions. After the instructions are synthesized, the corresponding answers are derived from the text.
@@ 3.3 Macro-foveate Level
At the broadest granularity, unsupervised text frequently employs figurative language and rhetorical devices (metaphor, hyperbole, contrastive foil, rhetorical questioning, and citation) that convey meaning beyond their literal content. A teacher LLM may overlook these devices without explicit prompting, missing the deeper communicative intent they carry. The macro-foveate level captures this holistic layer by identifying and extracting text segments that employ such writing techniques, then converting them into instructional form through transcription synthesis, ensuring that literary and rhetorical nuances are reflected in the synthesized instruction sets.
@@
This mechanism directs the teacher LLM to comprehend unsupervised text from a holistic perspective. We highlight and extract the text segments employing such writing techniques, which are designated as foveate segments.
@@
Based on the identified foveate segments, we employ transcription synthesis to convert each one into an instructional format. This process transforms declarative foveate segments into interrogative or imperative forms. Subsequently, corresponding answers are synthesized according to the content of the unsupervised text.
@@ 3.4 Re-synthesis Module
Due to limitations of the teacher LLM, not all instructions in a directly synthesized instruction set may be answerable from the source text. To improve fidelity, we employ a re-synthesis module that iteratively re-synthesizes these unanswerable instructions over multiple rounds, replacing them with instructions that can be fully answered based on the same text.
Single-Sample Reference Synthesis
During re-synthesis, we process one failed instruction at a time and randomly sample a different subset of successfully synthesized instructions as reference examples in each iteration, improving the success rate by exposing the teacher LLM to varied reference contexts.
@@
The outputs of LLMs are significantly influenced by hyperparameters such as temperature, top-p, and frequency penalty. To improve re-synthesis success rates, we tune these hyperparameters to define a high-creativity configuration, enabling the teacher LLM to synthesize more varied instructions based on reference samples.
Dataset Source # Test Examples Question Context SQuAD Crowdsourced Wikipedia 11639 HotpotQA Crowdsourced Wikipedia 2500 FilmWiki LLM Wikipedia 7398
@@
Datasets Methods Diversity Metrics SelfBLEU Div. Embedding Div. SQuAD Self-QA 0.593 0.838 Bonito 0.494 0.838 Wiki2023 0.550 0.842 Self-Foveate 0.665 0.851 Test Questions 0.695 0.840 HotpotQA Self-QA 0.463 0.823 Bonito 0.371 0.769 Wiki2023 0.554 0.822 Self-Foveate 0.607 0.835 Test Questions 0.634 0.786 FilmWiki Self-QA 0.406 0.687 Bonito 0.197 0.677 Wiki2023 0.341 0.664 Self-Foveate 0.563 0.706 Test Questions 0.316 0.618
@@
Model Settings GPT-4o mini DeepSeek-V3 SQuAD HotpotQA FilmWiki SQuAD HotpotQA FilmWiki Rec. Acc. Rec. Acc. Rec. Acc. Rec. Acc. Rec. Acc. Rec. Acc. Llama-3.1-8B None* 0.309 0.202 0.244 0.160 0.212 0.082 0.309 0.202 0.244 0.160 0.212 0.082 Self-QA 0.367 0.384 0.372 0.358 0.328 0.201 0.389 0.412 0.399 0.378 0.370 0.239 Wiki2023 0.327 0.361 0.338 0.322 0.333 0.235 0.342 0.370 0.340 0.328 0.349 0.244 Bonito* 0.386 0.405 0.360 0.372 0.219 0.153 0.386 0.405 0.360 0.372 0.219 0.153 Self-Foveate 0.484 0.490 0.507 0.486 0.512 0.367 0.481 0.491 0.525 0.501 0.548 0.397 Qwen2.5-7B None* 0.251 0.300 0.266 0.234 0.139 0.032 0.251 0.300 0.266 0.234 0.139 0.032 Self-QA 0.249 0.232 0.276 0.246 0.206 0.082 0.119 0.125 0.102 0.106 0.111 0.056 Wiki2023 0.215 0.221 0.135 0.112 0.192 0.093 0.170 0.083 0.197 0.203 0.202 0.136 Bonito* 0.143 0.109 0.212 0.199 0.168 0.098 0.143 0.109 0.212 0.199 0.168 0.098 Self-Foveate 0.408 0.414 0.372 0.329 0.283 0.140 0.388 0.389 0.342 0.331 0.261 0.140 Gemma-2-9B None* 0.224 0.121 0.175 0.078 0.211 0.099 0.224 0.121 0.175 0.078 0.221 0.099 Self-QA 0.383 0.409 0.408 0.389 0.429 0.315 0.402 0.435 0.424 0.408 0.509 0.386 Wiki2023 0.336 0.378 0.361 0.352 0.478 0.384 0.364 0.399 0.373 0.365 0.494 0.401 Bonito* 0.411 0.457 0.366 0.373 0.255 0.196 0.411 0.457 0.366 0.373 0.255 0.196 Self-Foveate 0.507 0.525 0.537 0.520 0.672 0.528 0.499 0.514 0.552 0.525 0.697 0.581
@@ 4 Experiment
Our experiments focus on three critical research questions: (Q1) How effective is Self-Foveate in enhancing instruction diversity? (Q2) How effective is Self-Foveate in enhancing instruction difficulty? (Q3) How effective is Self-Foveate in improving the model’s problem-solving capabilities during instruction fine-tuning?
@@ 4.1 Experimental Setup
We employ three independent datasets, including the training set of the FilmWiki dataset containing 2,385 unsupervised texts with corresponding question-answer pairs. Additionally, we sample two widely used benchmark QA datasets from the MRQA 2019 shared task (Fisch et al., 2019): SQuAD (Rajpurkar et al., 2016) (following Bonito (Nayak et al., 2024a)) and HotpotQA (Yang et al., 2018). To maintain comparable data scale and computational costs to the FilmWiki dataset, we extract 2,500 unsupervised texts with corresponding QA pairs from each dataset’s training split as substitutes for the complete collections. Further implementation details are provided in Table 2.
@@
We consider four baselines: zero-shot, Self-QA, Bonito, and Wiki2023. For the zero-shot baseline, we prompt models for evaluation without leveraging any unsupervised texts from the target task (None). The Self-QA baseline employs an unsupervised knowledge-guided method for extracting instruction-question-answer triples (Self-QA(Zhang and Yang, 2023). The Bonito baseline utilizes a 7B-sized specialized model to generate various types of questions from unsupervised text. We configure it to produce questions of the “question answering without choices” type and obtain a sufficient number of samples through multiple sampling iterations (Bonito). The Wiki2023 baseline implements a text-based QA pair extraction methodology (Wiki2023). All baselines use the same unsupervised texts as our method, as specified in Section 4.1.
@@
As described in section 3, we process unsupervised texts through Self-Foveate using GPT-4o mini and DeepSeek-V3 to synthesize instructional data. We implement two distinct hyperparameter configurations emphasizing high stability and high creativity, respectively, with detailed specifications in Appendix B.
@@
We select three state-of-the-art open-source foundation models as our base models prior to instruction tuning: Meta-Llama-3.1-8B (Grattafiori et al., 2024), Qwen2.5-7B (Qwen et al., 2025), and Gemma-2-9B (Team et al., 2024). These decoder-only language models employ next-word prediction objectives and were pretrained on trillions of tokens without any instruction-based fine-tuning.
4.2 Diversity Analysis
To explicitly evaluate the diversity of the instructions in a metric-driven manner, we follow established practices recommended by Zhu et al. (2018); Perez et al. (2022); Tevet and Berant (2021) and employ two metrics: SelfBLEU scores and Sentence-BERT embedding distances (Reimers and Gurevych, 2019a). These metrics capture different facets of diversity. SelfBLEU measures surface-level textual diversity, while embedding distances measure semantic diversity. For SelfBLEU, we compute average scores using $n$-grams for $n\in\{2,3,4,5\}$, following the approach suggested by Zhu et al. (2018). We use the implementation of the SelfBLEU metric by Alihosseini et al. (2019). Further details are available in Appendix G.
Table 3 demonstrates that Self-Foveate-synthesized instructions achieve substantial diversity improvements in both textual and semantic dimensions, attaining or even surpassing the diversity level of crowdsourced test questions through a low-cost automated process.
4.3 Difficulty Analysis
To investigate the difficulty of the synthesized instructions, we design prompts and employ GPT-4o to conduct a rigorous head-to-head comparison between the instructions synthesized by Self-Foveate and those synthesized by baseline methods. Specifically, for each unsupervised text in each dataset, we use two contrasting methods to synthesize an equal number of instructions. We then provide GPT-4o with a set of instructions synthesized by both methods under the same unsupervised text, ensuring that the instruction sets are anonymous and their relative positions are randomized. Subsequently, we record GPT-4o’s judgments and calculate the win rate for Self-Foveate. The results of the head-to-head comparison, as shown in Table 5, demonstrate the consistently higher difficulty of the instructions synthesized by Self-Foveate. Further details can be found in Appendix H.
Figure 3: Impact of instruction set scale on model fine-tuning performance for Self-Foveate and baselines: (a) Recall and (b) LLM Accuracy.
Dataset Baseline Self-Foveate WR. Win Lose SQuAD Self-QA 70.64% 29.36% Wiki2023 80.83% 19.17% Bonito 99.96% 00.04% HotpotQA Self-QA 89.52% 10.48% Wiki2023 91.17% 08.83% Bonito 100.00% 00.00% FilmWiki Self-QA 85.12% 14.88% Wiki2023 95.08% 04.92% Bonito 96.31% 03.69%
@@ 4.4 Problem-solving Capabilities
To evaluate the impact of instruction-tuning datasets on model performance in downstream tasks, we fine-tune open-source models and assess their performance on in-distribution problems. Table 4 illustrates the fine-tuning performance of open-source models using instruction sets synthesized by two teacher LLMs (GPT-4o mini and DeepSeek-V3), with equal quantities of instructions per unsupervised text. As answers in test problems often consist of phrase-level or sentence-fragment extractions from source texts while large language models (LLMs) tend to generate more comprehensive responses, we select both recall and LLM-evaluated accuracy as evaluation metrics. Higher recall indicates that generated answers are more likely to contain correct responses, while higher LLM-evaluated accuracy reflects stronger agreement between model-produced answers and ground truth labels, as judged by an LLM evaluator.
To control potential confounding effects from the number of instructions synthesized per unsupervised text and examine the agreement between these two metrics, we investigate the influence of the scale of instruction sets synthesized by Self-Foveate and baseline methods on model fine-tuning performance using the HotpotQA dataset, as shown in Figure 3. Additional details regarding downstream task evaluations are provided in Appendix I.
Table 4 demonstrates that all three base models achieve the best performance across downstream tasks in both metrics when fine-tuned with instructions synthesized via Self-Foveate, regardless of the teacher LLM used for synthesis. Notably, for Qwen2.5-7B, instruction data synthesized by baseline methods can lead to performance degradation compared to the base model without fine-tuning, further demonstrating the superiority of Self-Foveate in instruction synthesis. Our analysis in Figure 3 reveals that as the number of instructions synthesized by Self-Foveate per unsupervised text increases, downstream performance improves significantly, with an expanding gap compared to baseline methods. Figure 3 also confirms the consistency between recall and LLM-evaluated accuracy across varying instruction set scales.